OpenStudy (rainbow_dash):

Log(sub 2) 3 + log(sub 2) a = log(sub 2) 20

OpenStudy (anonymous):

\[\log_2(3)+\log_2(a)=\log_2(20)\]?

OpenStudy (anonymous):

start by combining the logs on the left using \[\log(x)+\log(y)=\log(xy)\] you get \[\log_2(3a)=\log(20)\]

OpenStudy (anonymous):

then since log is a one to one function, this means \[3a=20\] solve for \(a\)